OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundredth. 5x + 14 = 26

OpenStudy (smurfy14):

x=12/5

OpenStudy (anonymous):

5x=26-14 x= 12/5 x =2.4

OpenStudy (smurfy14):

in decimal form it is 2.40

OpenStudy (smurfy14):

nancy 2.4 is only rounding it to the tenth it is asking for the hundreth place

OpenStudy (anonymous):

yes hundreth is 2.40

OpenStudy (anonymous):

A. 1.45B. 1.47C. 1.54D. 1.58 Those are the answer choices. I don't understand how they would get any of those I got 2.4

OpenStudy (anonymous):

retype the problem

OpenStudy (anonymous):

5^x + 14 = 26

OpenStudy (smurfy14):

1.54

OpenStudy (anonymous):

Its 2.4 , there is nowhere you can round that to nearest hundred

OpenStudy (anonymous):

you said 5x +14 = 26 before

OpenStudy (anonymous):

the *round to the nearest hunderdth* is really weird

OpenStudy (anonymous):

Actually, this is the wrong question. Am I right?

OpenStudy (anonymous):

absolutely

OpenStudy (anonymous):

let's just leave.

OpenStudy (anonymous):

Which course are you upto OINH